# Institutskii per., 9, 141702, Dolgoprudny, Russia

### v2

@@ Abstract
We construct bi- and uni-vector deformations of 10d heterotic supergravity solutions with the GDFT approach. We construct a generalization of the “open/closed” map for this case and consider some examples of the deformed solutions, particularly for the F1 string solution.
@@ 2 Embedding of heterotic supergravity in GDFT
We consider a dimensionally reduced O(D,D+n) theory living in the space with coordinates $X^{M}=\left(x^{n},\tilde{x}_{n},y^{\alpha}\right)$, where $\left(x^{n},\tilde{x}_{n}\right)=\mathbb{X}$ are the external doubled space coordinates and $y^{\alpha}=\mathbb{Y}$ are the internal coordinates. After the Scherk-Schwarz compactification of the $\mathbb{Y}$-space, we get the action [44]
@@
$\left.+\left(S^{AB}-\eta^{AB}\right){F}_{A}{F}_{B}\right]\,$,
[…]
$\eta_{MN}=\left(\begin{array}[]{lll}0&\delta_{m}{}^{n}&0\\ \delta^{m}{}_{n}&0&0\\ 0&0&\delta_{\alpha\beta}\end{array}\right)\,.$ (2.2)
The fluxes in the reduced theory are given by
@@
Letters from the beginning of the Latin alphabet correspond to the tangent space, a vector in the tangent space having components $v^{A}=\left(v^{a}\,,v_{a}\,,v^{\bar{\alpha}}\right)$. Equations of motion can be obtained by varying with respect to the generalized dilaton $d$ and to the generalized vielbein $E_{A}{}^{M}$, taking into account the fact that it is an O(D,D+n) matrix, meaning that, $\delta E_{AB}\equiv E_{AM}\delta E_{B}{}^{M}=-E_{B}{}^{M}\delta E_{AM}$ is antisymmetric:
@@
where $Q_{ABC}=F_{ABC}+f_{ABC}\,$. The generalized vielbein can be Lorentz transformed to the form corresponding to the heterotic supergravity parameterization
@@
associated with the generalized metric defined as ${\mathcal{H}}_{MN}={E}^{A}{}_{M}{E}^{B}{}_{N}S_{AB}\,$,
$S_{AB}=\left(\begin{array}[]{lll}s_{ab}&0&0\\ 0&s^{ab}&0\\ 0&0&\delta_{\bar{\alpha}\bar{\beta}}\end{array}\right)\,$, (2.7)
@@
where $s_{ab}=$ diag(++)(-+...+) is the D-dimensional flat Minkowski metric, ${e}_{a}{}^{n}$ is the D-dimensional vielbein associated with $s_{ab}$ and ${c}_{mn}={b}_{mn}+\dfrac{1}{2}{A}_{m\gamma}{A}_{n}{}^{\gamma}\,$, and indices $\alpha,\beta$ are lowered and raised by $\delta_{\alpha\beta}$ and $\delta^{\alpha\beta}$. Finally, the invariant dilaton in terms of supergravity fields takes the following form
@@
The heterotic supergravity action can be obtained from (2.1) by substituting the fluxes $F_{ABC}$ using the parametrization of the generalized vielbein (2.6) and the invariant dilaton (2.9) and solving the section condition by only leaving the dependence on coordinates $x^{m}$ ($\frac{\partial}{\partial\tilde{x}_{n}}=\frac{\partial}{\partial y^{\alpha}}=0$):
@@ 3 Bi- and uni-vector deformations
To construct a deformation we follow the ideas of [27, 28, 29, 45, 46, 47]. We define a deformation as a simultaneous rotation of the generalized vielbein/metric and the structure constants [^1] We transform structure constants as they enter the action on equal footing with the fluxes $F_{MNK}$ and such transformation leaves the action invariant.
@@
with the O(D,D+n) matrix
@@
Let us now comment on the transformation of structure constants in (3.1). To remain in the heterotic theory after the deformation they should transform in such a way, that the structure constants $f_{\alpha\beta\gamma}$ in the action and equations of motion do not change and other non-trivial components like $\gamma^{m}{}_{\alpha}f^{\alpha\beta\gamma}$, $\gamma^{m}{}_{\alpha}\gamma^{n}{}_{\beta}f^{\alpha\beta\gamma}$, $\gamma^{m}{}_{\alpha}\gamma^{n}{}_{\beta}\gamma^{k}{}_{\gamma}f^{\alpha\beta\gamma}$ do not appear. Such terms enter the equations of motion and the action only as commutators $\gamma^{m},A_{n}]$ and $\gamma^{m},\gamma^{n}]$, where $\gamma^{m}=\gamma^{m}{}_{\alpha}\mathfrak{t}^{\alpha}$ and $A_{m}=A_{m\alpha}\mathfrak{t}^{\alpha}$. Hence for the deformation (3.2) to work, we need $\gamma^{m}{}_{\alpha}$ to satisfy
@@ 5 Conclusion and discussion
The third interesting direction is to investigate the preservation of integrability and kappa-symmetry for the two-dimensional sigma model on the deformed heterotic backgrounds, which also has direct holographic implications for LSTs. For these purposes quite useful approach is gauged double sigma model [68]. Furthermore, the recent work [69] has explored integrable heterotic deformations of the principal chiral model that turn on gauge fields. A natural and intriguing question is whether this deformation can be interpreted as a uni-vector heterotic deformation.

### v3

@@ Abstract
We construct bi- and uni-vector deformations of 10d heterotic supergravity solutions with the gauged double field theory approach. We construct a generalization of the “open/closed” map for this case and consider some examples of the deformed solutions, particularly for the F1 string solution.
@@ 2 Embedding of heterotic supergravity in GDFT
We consider a dimensionally reduced O(D,D+n) theory living in the space with coordinates $X^{M}=\left(x^{m},\tilde{x}_{m},y^{\alpha}\right)$, where $\left(x^{m},\tilde{x}_{m}\right)=\mathbb{X}$ are the external doubled space coordinates and $y^{\alpha}=\mathbb{Y}$ are the internal coordinates. After the Scherk-Schwarz compactification of the $\mathbb{Y}$-space, we get the action [44]
@@
$\left.+\left(S^{AB}-\eta^{AB}\right){F}_{A}{F}_{B}\right]\,.$
In the above action nothing depends on $y^{\alpha}$, and $v$ is a constant appearing from the integration over the reduced coordinates $\mathbb{Y}$ (see [44] for details), $d$ is O(D,D+n) invariant dilaton, $S^{AB}$ is inverse to flat generalized metric $S_{AB}$, and $\eta^{AB}$ is inverse to O(D,D+n) invariant metric $\eta_{AB}$ :
$S_{AB}=\left(\begin{array}[]{lll}s_{ab}&0&0\\ 0&s^{ab}&0\\ 0&0&\delta_{\bar{\alpha}\bar{\beta}}\end{array}\right)\,,\quad\eta_{AB}=\left(\begin{array}[]{lll}0&\delta_{a}{}^{b}&0\\ \delta^{a}{}_{b}&0&0\\ 0&0&\delta_{\bar{\alpha}\bar{\beta}}\end{array}\right)\,$, (2.2)
where $s_{ab}=$ diag(++)(-+...+) is the D-dimensional flat Minkowski metric. Flat indices are transformed to the curved ones by contraction with generalized vielbein $E_{A}{}^{M}\in$ O(D,D+n) and are lowered and raised by O(D,D+n) invariant metric $\eta$. Curved generalized metric is defined as ${\mathcal{H}}_{MN}={E}^{A}{}_{M}{E}^{B}{}_{N}S_{AB}$ and fluxes with the flat indices are related to the curved ones as $f_{ABC}=E_{A}{}^{M}E_{B}{}^{N}E_{C}{}^{K}f_{MNK}$, $F_{ABC}=E_{A}{}^{M}E_{B}{}^{N}E_{C}{}^{K}F_{MNK}$, where
@@
[…]
@@
where $Q_{ABC}=F_{ABC}+f_{ABC}\,$. The generalized vielbein can be Lorentz transformed to the form corresponding to the heterotic supergravity parameterization [^1] Other parameterizations of GDFT can be used to describe non-geometric backgrounds.
@@
associated with the generalized metric ${\mathcal{H}}_{MN}={E}^{A}{}_{M}{E}^{B}{}_{N}S_{AB}\,$,
@@
where ${e}_{a}{}^{n}$ is the D-dimensional vielbein associated with $s_{ab}$ and ${c}_{mn}={b}_{mn}+\dfrac{1}{2}{A}_{m\gamma}{A}_{n}{}^{\gamma}\,$, and indices $\alpha,\beta$ are lowered and raised by $\delta_{\alpha\beta}$ and $\delta^{\alpha\beta}$. Finally, the invariant dilaton in terms of supergravity fields takes the following form
@@
The heterotic supergravity action can be obtained from (2.1) by substituting the fluxes $F_{ABC}$ using the parametrization of the generalized vielbein (2.6) and the invariant dilaton (2.8) and solving the section constraint by only leaving the dependence on coordinates $x^{m}$ ($\frac{\partial}{\partial\tilde{x}_{n}}=0$):
@@ 3 Bi- and uni-vector deformations
To construct a deformation we follow the ideas of [27, 28, 29, 45, 46, 47]. We define a deformation as a simultaneous rotation of the generalized vielbein/metric and the structure constants
@@
with the O(D,D+n) matrix $O_{M}{}^{N}$ defined through the corresponding generators $T_{mn}$, $T_{m\alpha}$ of o(D,D+n)
@@
Let us now comment on the transformation of $f_{MNK}$ in (3.1). The only non-zero components of $f^{MNK}$ are $f^{\alpha\beta\gamma}\,$, which are the fixed constants in our theory. However, their transformation yields additional components $f^{\prime m\beta\gamma}=\gamma^{m}{}_{\alpha}f^{\alpha\beta\gamma}$, $f^{\prime mn\gamma}=\gamma^{m}{}_{\alpha}\gamma^{n}{}_{\beta}f^{\alpha\beta\gamma}$, $f^{\prime mnk}=\gamma^{m}{}_{\alpha}\gamma^{n}{}_{\beta}\gamma^{k}{}_{\gamma}f^{\alpha\beta\gamma}$. To remain in the heterotic theory the contribution of these components should vanish in the equations of motion. Such terms enter the equations of motion and the action only as commutators $\gamma^{m},A_{n}]$ and $\gamma^{m},\gamma^{n}]$, where $\gamma^{m}=\gamma^{m}{}_{\alpha}\mathfrak{t}^{\alpha}$ and $A_{m}=A_{m\alpha}\mathfrak{t}^{\alpha}$. Hence for the deformation (3.2) to work, we need $\gamma^{m}{}_{\alpha}$ to satisfy
@@ 5 Conclusion and discussion
The third interesting direction is to investigate the preservation of integrability and kappa-symmetry for the two-dimensional sigma model on the deformed heterotic backgrounds, which also has direct holographic implications for LSTs. For these purposes quite useful approach is gauged double sigma model [68]. Furthermore, the recent work [69] has explored integrable heterotic deformations of the principal chiral model that turn on gauge fields. A natural and intriguing question is whether this deformation can be interpreted as a uni-vector heterotic deformation. Also, recent work [70] has taken the first steps toward understanding integrability for poly-vector deformed backgrounds.